find the reciprocal and express it in scientific notation.
6.25*10^-3
Please help

Answer by Alan3354(30993)

find the reciprocal and express it in scientific notation.
6.25*10^-3
-----------
1/6.25*10^-3 = (1/6.25)*10^3 or (1/6.25)E3
= 0.16E3
= 1.6E2
-------
= 1.6*10^2 if you prefer inconvenience